40 Coulombs of charge are separated by 6 km in a thundercloud.  How large is the electric field half way between the charges?  This gives an idea of large electric fields you might experience
Answer #1

An electric field is a vector field that associates to each point in space the Coulomb force experienced by a unit electric charge. \(^{\text {[1] Electric fields converge and diverge at electric charges and they can be induced by }}\) time-varying magnetic fields. The electric field combines with the magnetic field to form the electromagnetic field.

Electric field \(\mathrm{E}=\mathrm{k} \mathrm{q} / \mathrm{r}^{2}\)

$$ \begin{array}{l} r=\text { separation }=3 \mathrm{~km}=3 * 10^{3} \mathrm{~m} \\ q=40 \mathrm{C} \end{array} $$

Electric field \(=9 * 10^{9 *} 40 /\left(6^{*} 10^{3}\right)^{2}\)

$$ =10^{4} \mathrm{~N} / \mathrm{C} $$

when it is half way , then \(r=3^{*} 10^{3} \mathrm{~m}\)

Electric field \(=9 * 10^{9 *} 40 /\left(3^{*} 10^{3}\right)^{2}\)

\(=4 * 10^{4} \mathrm{~N} / \mathrm{C}\)

It is nearly 4 times greater than the original field between them. It's too large.

answered by: Firster